Please help with these questions
Limiting Reactants and Theoretical Yield Calculations 15 30 20g) lf 112g of N2 reacts completely with 30.0g H2. l) what is the limiting reactant? 4 mdes 15 mde 3 5 moes 2. 2) What is the theoretical yield? 30 Hz x N limitiry 3) How much of the reagent in excess is left over? As
Answer #1

For the given reaction,

moles of N2 = 112 g/28 g/mol = 4 mols

moles of H2 = 30.2 g/2 = 15.1 mols

If all of N2 is consumed, we would need = 12 mols

1) Since moles of H2 available is greater than needed, N2 is the limiting reactant

2) theoretical yield of NH3 = 4 mol x 2 x 17 g/mol = 136 g

3) Excess reagent left at the end = (15.1 - 12) mols x 2 = 6.2 g